# Diffraction Aided Wireless Positioning

### v1

@@ III-A Simplified Model of a Building
$\hat{\phi}^{\prime}$ $\frac{(z_{e}^{l}-z_{a})\hat{y}-(y_{a})\hat{z}}{\sqrt{(y_{a})^{2}+(z_{e}^{l}-z_{a})^{2}}}$ $\frac{(z_{e}^{u}-z_{a})\hat{y}+(y_{a})\hat{z}}{\sqrt{(y_{a})^{2}+(z_{e}^{u}-z_{a})^{2})}}$
$\hat{\bm{\beta}}^{\prime}_{0}$ $\frac{-(y_{a}^{2}+(z_{e}^{l}-z_{a})^{2})\hat{x}-(y_{a})(q_{x}-x_{a})\hat{y}+(z_{e}^{l}-z_{a})(q_{x}-x_{a})\hat{z}}{\sqrt{(y_{a})^{2}+(z_{e}^{l}-z_{a})^{2}}\sqrt{(q_{x}-x_{a})^{2}+(y_{a})^{2}+(z_{e}^{l}-z_{a})^{2}}}$ $\frac{(y_{a}^{2}+(z_{e}^{u}-z_{a})^{2})\hat{x}+(q_{x}-x_{a})\hat{y}-(q_{x}-x_{a})(z_{e}^{u}-z_{a})\hat{z}}{\sqrt{(y_{a})^{2}+(z_{e}^{u}-z_{a})^{2}}\sqrt{(q_{x}-x_{a})^{2}+(y_{a})^{2}+(z_{e}^{u}-z_{a})^{2}}}$
$\hat{\bm{\phi}}$ $\frac{(0.5w)\hat{y}-(y_{n})\hat{z}}{\sqrt{(y_{n})^{2}+(0.5w)^{2}}}$ $\frac{(0.5w)\hat{y}+(y_{n})\hat{z}}{\sqrt{(y_{n})^{2}+(0.5w)^{2}}}$
$\hat{\bm{\beta}_{0}}$ $\frac{(y_{n}^{2}+(0.5w)^{2})\hat{x}-(y_{n})(x_{n}-q_{x})\hat{y}+(0.5w)(x_{n}-q_{x})\hat{z}}{\sqrt{(y_{n})^{2}+(0.5w)^{2}}\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}}$ $\frac{-(y_{n}^{2}+(0.5w)^{2})\hat{x}+(y_{n})(x_{n}-q_{x})\hat{y}-(0.5w)(x_{n}-q_{x})\hat{z}}{\sqrt{(y_{n})^{2}+(0.5w)^{2}}\sqrt{(x_{n}-q_{x})^{2}+(y_{n})^{2}+(0.5w)^{2}}}$
Assume the anchor transmits a linearly polarized electric field $\bm{E}_{0}=(E_{0x},E_{0y},E_{0z})^{T}$, for each edge, the incident field at the diffraction point $Q_{e}$ lying on the respective edge is expressed as a vector with components along orthogonal vectors $(\hat{\bm{\beta}}_{0}^{\prime})$ and $(\hat{\bm{\phi}}^{\prime})$. These orthogonal vectors are calculated using eq. (8) for both edges. Similarly, the diffracted electric field for each edge is obtained along two orthogonal vectors $(\hat{\bm{\beta}}_{0})$ and $(\hat{\bm{\phi}})$ using eq. (7). Let the incident field orthogonal vectors be $\hat{\bm{\beta}}_{0}^{\prime}=(\hat{\bm{\beta}}_{0x}^{\prime},\hat{\bm{\beta}}_{0y}^{\prime},\hat{\bm{\beta}}_{0z}^{\prime})$, $\hat{\bm{\phi}}^{\prime}=(\hat{\bm{\phi}_{x}}^{\prime},\hat{\bm{\phi}_{y}}^{\prime},\hat{\bm{\phi}_{z}}^{\prime})$ and the diffracted field orthogonal vectors be $\hat{\bm{\beta}}_{0}=(\hat{\bm{\beta}}_{0x},\hat{\bm{\beta}}_{0y},\hat{\bm{\beta}}_{0z})$, $\hat{\bm{\phi}}=(\hat{\bm{\phi}_{x}},\hat{\bm{\phi}_{y}},\hat{\bm{\phi}_{z}})$ where the subscript ‘$x$’,‘$y$’ and ‘$z$’ represents the component along the X,Y and Z axes respectively.
@@
$\begin{split}\bm{E}^{d}&=\left[\left(E_{0x}\hat{\bm{\beta}}_{0x}^{\prime}+E_{0y}\hat{\bm{\beta}}_{0y}^{\prime}+E_{0z}\hat{\bm{\beta}}_{0z}^{\prime}\right)\right.\\ &\left.\left(\hat{\beta}_{0x}\hat{x}+\hat{\beta}_{0y}\hat{y}+\hat{\beta}_{0z}\hat{z}\right)D_{s}+\left(E_{0y}\hat{\bm{\phi}_{y}}^{\prime}+E_{0z}\hat{\bm{\phi}_{z}}^{\prime}\right)\right.\\ &\left.\left(0\hat{x}+\hat{\bm{\phi}_{y}}\hat{y}+\hat{\bm{\phi}_{z}}\hat{z}\right)D_{h}\right]\frac{\mathrm{e}^{-jk(|s^{{}^{\prime}}|+|s|)}}{|s^{{}^{\prime}}|\sqrt{|s|}}.\end{split}$ (13)
@@
The required orthogonal vectors for the incident and diffracted electric field for the upper and lower edge diffraction MPCs are presented in Table II. Note that $\hat{\bm{\phi}_{x}}^{\prime}=0$ and $\hat{\bm{\phi}_{x}}=0$ for both edges. Hence, the diffraction electric field for each edge can be obtained using Definition 2 and the vectors for each edge. ∎
@@ III-B Relative power of the two diffraction MPCs
In this section, we examine the effect of controlling the transmit side electric field polarization on the relative power of the two diffraction paths. First, we derive expressions for the soft and hard diffraction coefficients for the two diffraction MPCs as a function of the elevation angle $\theta$ as in Fig. 1(b). From the definition of the diffraction coefficients in eq. (9) we first need to obtain angles $\psi^{\prime}$ and $\psi$ using eq. (11). We derive $s_{t}^{\prime}$ and $s_{t}$ for both edges using eq. (10) and the values in Table I, Table II.
@@
To calculate $\psi$ and $\psi^{\prime}$ for both edges, we assume
@@
  • Define angle $\theta=\cos^{-1}{\left(\frac{z_{e}^{l}-z_{a}}{\sqrt{(y_{a})^{2}+(z_{e}^{l}-z_{a})^{2}}}\right)}\approx\cos^{-1}{\left(\frac{z_{e}^{u}-z_{a}}{\sqrt{(y_{a})^{2}+(z_{e}^{u}-z_{a})^{2}}}\right)}$ as shown in Fig. 1(b).
@@
  • The diffraction MPCs arrive at the node location approximately parallel to the floor. In other words, the elevation AoA for the diffraction MPCs at the receiver will be approximately  [^1] Our observation about the AoA of the diffraction MPCs arriving parallel to the floor also matches the observation using Raytracing software in Appendix -D.$90^{\circ}$. This result is obtained rigorously by substituting statement 1 of Assumption 2 into the expression for the angle $\psi$ presented in Table III to obtain $\psi\approx\frac{3\pi}{2}$. The angle $\psi$ is shown in Fig. 1(a).
  • The path length of the two diffraction MPCs generated by the upper and lower window edges is approximately the same. This is obtained by applying the assumption to the IPL and OPL defined in corollary 1 for both upper and lower edge diffraction MPCs.
TABLE III: $s^{\prime}_{t}$, $s_{t}$, $\hat{\psi}^{\prime}$, $\hat{\psi}$ for upper and lower edges.
@@
$s^{\prime}_{t}$ $\frac{-y_{a}\hat{y}+(z_{e}^{l}-z_{a})\hat{z}}{\sqrt{(y_{a})^{2}+(z_{e}^{l}-z_{a})^{2}}}$ $\frac{-y_{a}\hat{y}+(z_{e}^{u}-z_{a})\hat{z}}{\sqrt{(y_{a})^{2}+(z_{e}^{u}-z_{a})^{2}}}$
$s_{t}$ $\frac{y_{n}\hat{y}+(0.5w)\hat{z}}{\sqrt{(y_{n})^{2}+(0.5w)^{2}}}$ $\frac{y_{n}\hat{y}-(0.5w)\hat{z}}{\sqrt{(y_{n})^{2}+(0.5w)^{2}}}$
$\psi^{\prime}$ $\arccos{\left(\frac{(z_{e}^{l}-z_{a})}{\sqrt{(y_{a})^{2}+(z_{e}^{l}-z_{a})^{2}}}\right)}=\theta$ $\pi-\arccos{\left(\frac{(z_{e}^{u}-z_{a})}{\sqrt{(y_{a})^{2}+(z_{e}^{u}-z_{a})^{2}}}\right)}\approx\pi-\theta$
@@
The second statement in Assumption 2 leads to relating angles $\psi^{\prime}$ for both the diffraction paths with the UAV anchor elevation angle $\theta$ shown in 1(b). The expressions for $\psi^{\prime}$ for both diffraction MPCs are presented Table III. Observe from Fig. 1(a), for each diffraction MPC, $\psi^{\prime}$ is the angle the incident diffraction ray makes with the half-plane, and this suggests that the electric field strength depends on the UAV anchor elevation angle. To investigate the dependence of the electric field strength on the elevation angle $\theta$, we derive approximate expressions for the diffraction coefficients.
@@
The soft ($D_{s}$) and hard ($D_{h}$) diffraction coefficients for the two diffraction MPCs are presented in Table IV. Note the superscripts ‘u’ and ‘l’ refer to the upper and lower diffraction MPCs respectively.
@@
Soft Diff. Coeff. $D_{s}^{l}=A\left(\frac{-2\sqrt{1-\cos{\theta}}}{\cos{\theta}}\right)$ $D_{s}^{u}=A\left(\frac{2\sqrt{1+\cos{\theta}}}{\cos{\theta}}\right)$
Hard Diff. Coeff. $D_{h}^{l}=A\left(\frac{2\sqrt{1+\cos{\theta}}}{\cos{\theta}}\right)$ $D_{h}^{u}=A\left(\frac{-2\sqrt{1-\cos{\theta}}}{\cos{\theta}}\right)$
@@
Please refer to Appendix -C. ∎
Figure 2: Ratio of Power in the diffraction MPC for $\theta\in[0,90]$ for various node locations compared with the approximation in eq. (LABEL:eq_power_ratio).
In Fig. 2, we plot the approximate power ratio expression of the two diffraction MPCs in eq. (LABEL:eq_power_ratio) as a function of the elevation angle $\theta$ to the anchor. This is compared with the exact power ratio which depends on the node position and shows good agreement. For $\theta<60$°, i.e., anchor positions primarily in NLOS with the node, the diffraction MPC power ratio is atleast 5dB. Therefore, we can conclude that the upper edge diffraction MPC dominates over the lower edge diffraction MPC leading to a single dominating diffraction MPC.
@@ IV Indoor Positioning
Here $a_{rx,k}(\theta_{rx,k})$ and $a_{tx,k}(\theta_{tx,k})$ are the respective receiver and transmitter array response vectors for the $k^{th}$ MPC formed between the transmitter and the receiver. Each MPC has four properties: (a) complex gain: $h_{k}$, (b) Angle-of-Arrival (AoA): $\theta_{rx,k}$, (c) Angle-of-Departure(AoD): $\theta_{tx,k}$ and the propagation delay: $\tau_{k}$. We assume the transmitter and receiver antenna gains only vary across elevation, thus the path gain is only affected by the elevation AoA and AoD of the MPCs. All four parameters for each MPC are ultimately affected by objects in the environment and the frequency of operation through the three main mechanisms: (a) reflection from surfaces, (b) diffraction from edges, and (c) transmission through various building materials. Both the building materials and the geometry of the walls can attenuate the MPCs by changing the number of significant reflections, diffractions, and transmissions each path can undergo before it reaches the receiver. We conducted a statistical study on the MPCs generated for an O2I signal propagation scenario relevant to our positioning system. The key results used for further analysis are shown below, with the details of the study presented in Appendix -D.
@@
  • The diffraction MPCs arrive at the receiver nearly parallel to the ground, i.e., the receive side AoA elevation angle is approximately $90^{\circ}$.
IV-A TOF range measurements
To achieve our ultimate goal of 3D positioning, we are interested in estimating the path delay associated with the diffraction MPCs. Based on the previous section, this is ultimately equivalent to estimating the propagation delay of the first arriving path between the outside anchor and the indoor node. The signal’s propagation delay along the first arriving path will correspond to one of the two diffraction paths and can be estimated using the received signal measurements. Assume the anchor and node are  [^3] This is not a strict requirement. We could also use Round Trip Time (RTT) based systems as specified in 3GPP and WiFi standardsperfectly time synchronized. Now, we transmit a pulse of power $E_{p}$ and pulse shape $p(t)$ of duration $T_{p}$ at the transmitter. Convolving the transmit pulse with the channel defined in eq. (15) we obtain the received signal in the presence of Gaussian noise as

### v2

@@ III-A Simplified Model of a Building
ϕ^superscriptbold-^bold-italic-$\bm{\hat{\phi}^{\prime}}$ $\frac{(z_{e}^{l}-z_{a})\bm{\hat{y}}-(y_{a})\bm{\hat{z}}}{\sqrt{(y_{a})^{2}+(z_{e}^{l}-z_{a})^{2}}}$ $\frac{(z_{e}^{u}-z_{a})\bm{\hat{y}}+(y_{a})\bm{\hat{z}}}{\sqrt{(y_{a})^{2}+(z_{e}^{u}-z_{a})^{2}}}$
$\bm{\hat{\bm{\beta}}^{\prime}_{0}}$ $\frac{-(y_{a}^{2}+(z_{e}^{l}-z_{a})^{2})\bm{\hat{x}}-(y_{a})(q_{x}^{l}-x_{a})\bm{\hat{y}}+(z_{e}^{l}-z_{a})(q_{x}^{l}-x_{a})\bm{\hat{z}}}{\sqrt{(y_{a})^{2}+(z_{e}^{l}-z_{a})^{2}}\sqrt{(q_{x}^{l}-x_{a})^{2}+(y_{a})^{2}+(z_{e}^{l}-z_{a})^{2}}}$ $\frac{(y_{a}^{2}+(z_{e}^{u}-z_{a})^{2})\bm{\hat{x}}+(y_{a})(q_{x}^{u}-x_{a})\bm{\hat{y}}-(q_{x}^{u}-x_{a})(z_{e}^{u}-z_{a})\bm{\hat{z}}}{\sqrt{(y_{a})^{2}+(z_{e}^{u}-z_{a})^{2}}\sqrt{(q_{x}^{u}-x_{a})^{2}+(y_{a})^{2}+(z_{e}^{u}-z_{a})^{2}}}$
$\bm{\hat{\bm{\phi}}}$ $\frac{(0.5w)\bm{\hat{y}}-(y_{n})\bm{\hat{z}}}{\sqrt{(y_{n})^{2}+(0.5w)^{2}}}$ $\frac{(0.5w)\bm{\hat{y}}+(y_{n})\bm{\hat{z}}}{\sqrt{(y_{n})^{2}+(0.5w)^{2}}}$
$\bm{\hat{\bm{\beta}_{0}}}$ $\frac{(y_{n}^{2}+(0.5w)^{2})\bm{\hat{x}}-(y_{n})(x_{n}-q_{x}^{l})\bm{\hat{y}}+(0.5w)(x_{n}-q_{x}^{l})\bm{\hat{z}}}{\sqrt{(y_{n})^{2}+(0.5w)^{2}}\sqrt{(x_{n}-q_{x}^{l})^{2}+(y_{n})^{2}+(0.5w)^{2}}}$ $\frac{-(y_{n}^{2}+(0.5w)^{2})\bm{\hat{x}}+(y_{n})(x_{n}-q_{x}^{u})\bm{\hat{y}}-(0.5w)(x_{n}-q_{x}^{u})\bm{\hat{z}}}{\sqrt{(y_{n})^{2}+(0.5w)^{2}}\sqrt{(x_{n}-q_{x}^{u})^{2}+(y_{n})^{2}+(0.5w)^{2}}}$
Assume that the anchor transmits a linearly polarized electric field $\bm{E_{0}}=[E_{0x},E_{0y},E_{0z}]^{T}$, for each edge, the incident field at the diffraction point $\bm{Q_{e}}$ lying on the respective edge is expressed as a vector with components along orthogonal vectors $\bm{\hat{\bm{\beta}}_{0}^{\prime}}$ and ϕ^superscriptbold-^bold-italic-$\bm{\hat{\bm{\phi}}^{\prime}}$. These orthogonal vectors are calculated using eq. (8) for both edges. Similarly, the diffracted electric field for each edge is obtained along two orthogonal vectors $\bm{\hat{\bm{\beta}}_{0}}$ and $\bm{\hat{\bm{\phi}}}$ using eq. (7). Let the incident field orthogonal vectors be $\bm{\hat{\bm{\beta}}_{0}^{\prime}}=[\hat{\bm{\beta}}_{0x}^{\prime},\hat{\bm{\beta}}_{0y}^{\prime},\hat{\bm{\beta}}_{0z}^{\prime}]^{T}$, $\bm{\hat{\bm{\phi}}^{\prime}}=[\hat{\bm{\phi}_{x}}^{\prime},\hat{\bm{\phi}_{y}}^{\prime},\hat{\bm{\phi}_{z}}^{\prime}]^{T}$ and the diffracted field orthogonal vectors be $\bm{\hat{\bm{\beta}}_{0}}=[\hat{\bm{\beta}}_{0x},\hat{\bm{\beta}}_{0y},\hat{\bm{\beta}}_{0z}]^{T}$, $\bm{\hat{\bm{\phi}}}=[\hat{\bm{\phi}_{x}},\hat{\bm{\phi}_{y}},\hat{\bm{\phi}_{z}}]^{T}$ where the subscript ‘$x$’,‘$y$’ and ‘$z$’ represents the component along the X,Y and Z axes, respectively.
@@
$\begin{split}\bm{E^{d}}&=\left[\left(E_{0x}\hat{\beta}_{0x}^{\prime}+E_{0y}\hat{\beta}_{0y}^{\prime}+E_{0z}\hat{\beta}_{0z}^{\prime}\right)\right.\\ &\left.\left(\hat{\beta}_{0x}\bm{\hat{x}}+\hat{\beta}_{0y}\bm{\hat{y}}+\hat{\beta}_{0z}\bm{\hat{z}}\right)D_{s}+\left(E_{0y}\hat{\phi_{y}}^{\prime}+E_{0z}\hat{\phi}_{z}^{\prime}\right)\right.\\ &\left.\left(0\bm{\hat{x}}+\hat{\bm{\phi}_{y}}\bm{\hat{y}}+\hat{\bm{\phi}_{z}}\bm{\hat{z}}\right)D_{h}\right]\frac{\mathrm{e}^{-jk(|\bm{s^{{}^{\prime}}}|+|\bm{s}|)}}{|\bm{s^{{}^{\prime}}}|\sqrt{|\bm{s}|}}.\end{split}$ (13)
@@
The required orthogonal vectors for the incident and diffracted electric field for the upper and lower edge diffraction MPCs are presented in Table II. Note that the x-component of the vector ϕ^bold-^superscriptbold-italic-$\bm{\hat{\phi^{\prime}}}$, $\hat{\phi_{x}}^{\prime}=0$ and the x-component of the vector $\bm{\hat{\phi}}$, $\hat{\phi_{x}}=0$ for both edges. Hence, the diffraction electric field for each edge can be obtained using definition 2 and the vectors for each edge. ∎
@@ III-B Relative power of the two diffraction MPCs
In this section, we examine the effect of controlling the polarization of the transmit side electric field on the relative power of the two diffraction paths. First, we derive expressions for the soft and hard diffraction coefficients for the two diffraction MPCs as a function of the elevation angle $\theta$ as in Fig. 1(b). From the definition of the diffraction coefficients in eq. (9) we first need to obtain the angles $\psi^{\prime}$ and $\psi$ using eq. (11). We derive $\bm{s_{t}^{\prime}}$ and $\bm{s_{t}}$ for both edges using eq. (10) and the values in Table I, Table II.
@@
We assume
@@
  • Define angle $\theta\approx\cos^{-1}{\left(\frac{z_{e}^{l}-z_{a}}{\sqrt{(y_{a})^{2}+(z_{e}^{l}-z_{a})^{2}}}\right)}\approx\cos^{-1}{\left(\frac{z_{e}^{u}-z_{a}}{\sqrt{(y_{a})^{2}+(z_{e}^{u}-z_{a})^{2}}}\right)}$ as the elevation angle of the UAV-anchor from the perspective of the building floor.
Note, the above assumption is only used to derive insights and is relaxed in the numerical results in Fig. 3.
@@
  • The diffraction MPCs arrive at the node location approximately parallel to the floor. In other words, the elevation AoA for the diffraction MPCs at the receiver will be approximately  [^1] The AoA is defined as the vertical angle between the MPC propagation direction and the z-axis.Our observation about the AoA of the diffraction MPCs arriving parallel to the floor also matches the observation using Raytracing software in Appendix -D.$90^{\circ}$. This result is obtained rigorously by substituting statement 1 of Assumption 2 into the expression for the angle $\psi$ presented in Table III to obtain $\psi\approx\frac{3\pi}{2}$. The angle $\psi$ is shown in Fig. 1(a).
  • The path lengths of the two diffraction MPCs generated by the upper and lower window edges are approximately the same. This is obtained by applying the assumption to the IPL and OPL defined in Corollary1 for both upper and lower edge diffraction MPCs.
TABLE III: $\bm{s^{\prime}_{t}}$, $\bm{s_{t}}$, $\hat{\psi}^{\prime}$, $\hat{\psi}$ for upper and lower edges.
@@
$\bm{s^{\prime}_{t}}$ $\frac{-y_{a}\bm{\hat{y}}+(z_{e}^{l}-z_{a})\bm{\hat{z}}}{\sqrt{(y_{a})^{2}+(z_{e}^{l}-z_{a})^{2}}}$ $\frac{-y_{a}\bm{\hat{y}}+(z_{e}^{u}-z_{a})\bm{\hat{z}}}{\sqrt{(y_{a})^{2}+(z_{e}^{u}-z_{a})^{2}}}$
$\bm{s_{t}}$ $\frac{y_{n}\bm{\hat{y}}+(0.5w)\bm{\hat{z}}}{\sqrt{(y_{n})^{2}+(0.5w)^{2}}}$ $\frac{y_{n}\bm{\hat{y}}-(0.5w)\bm{\hat{z}}}{\sqrt{(y_{n})^{2}+(0.5w)^{2}}}$
$\psi^{\prime}$ $\arccos{\left(\frac{(z_{e}^{l}-z_{a})}{\sqrt{(y_{a})^{2}+(z_{e}^{l}-z_{a})^{2}}}\right)}\approx\theta$ $\pi-\arccos{\left(\frac{(z_{e}^{u}-z_{a})}{\sqrt{(y_{a})^{2}+(z_{e}^{u}-z_{a})^{2}}}\right)}\approx\pi-\theta$
@@
The second statement in Assumption 2 leads to relating angles $\psi^{\prime}$ for both the diffraction paths with the UAV anchor elevation angle $\theta$ shown in Fig.1(b). The expressions for $\psi^{\prime}$ for both diffraction MPCs are presented in Table III. Observe from Fig. 1(a), for each diffraction MPC, $\psi^{\prime}$ is the angle the incident diffraction ray makes with the half-plane, and this suggests that the electric field strength depends on the UAV anchor elevation angle. To investigate the dependence of the electric field strength on the elevation angle $\theta$, we derive approximate expressions for the diffraction coefficients.
@@
The soft ($D_{s}$) and hard ($D_{h}$) diffraction coefficients for the two diffraction MPCs are presented in Table IV. Note that the superscripts ‘u’ and ‘l’ refer to the upper and lower diffraction MPCs, respectively.
@@
Soft Diff. Coeff. $D_{s}^{l}\approx A\left(\frac{-2\sqrt{1-\cos{\theta}}}{\cos{\theta}}\right)$ $D_{s}^{u}\approx A\left(\frac{2\sqrt{1+\cos{\theta}}}{\cos{\theta}}\right)$
Hard Diff. Coeff. $D_{h}^{l}\approx A\left(\frac{2\sqrt{1+\cos{\theta}}}{\cos{\theta}}\right)$ $D_{h}^{u}\approx A\left(\frac{-2\sqrt{1-\cos{\theta}}}{\cos{\theta}}\right)$
@@
Refer to Appendix -C. ∎
Figure 3: Ratio of Power $\frac{P^{u}}{P_{l}}$ in the diffraction MPC for $\theta\in[0,90]$ for various node locations without Assumption 2 compared with the approximation in eq. (LABEL:eq_power_ratio) with Assumption 2. The elevation angle decreases as we go towards higher floors above the UAV anchor.
In Fig. 3, we plot the approximate power ratio expression of the two diffraction MPCs in eq. (LABEL:eq_power_ratio) as a function of the elevation angle $\theta$ to the anchor. This is compared with the exact power ratio which is calculated without Assumption 2 and depends on the node position within the building and shows good agreement. Note, $\theta=90$° represents the building floor at the same height as the UAV-anchor. Now, as we go towards higher floors, $\theta$ decreases and for $\theta<60$°, i.e., for building floors located above the height of the UAV-anchor, the diffraction MPC power ratio is at least 5dB. Therefore, we can conclude that for the locations of the nodes on the upper floors of the building, the diffraction MPC of the upper edge dominates the diffraction MPC of the lower edge, leading to a single dominating diffraction MPC.
@@ IV Indoor Positioning
Here, $a_{rx,k}(\theta_{rx,k})$ and $a_{tx,k}(\theta_{tx,k})$ are the respective gains of the receiver and transmitter antenna for the $k^{th}$ MPC formed between the transmitter and the receiver. Each MPC has four properties: (a) complex gain: $h_{k}$, (b) Angle-of-Arrival (AoA): $\theta_{rx,k}$, (c) Angle-of-Departure(AoD): $\theta_{tx,k}$ and (d) the propagation delay: $\tau_{k}$. We assume that the transmitter and receiver antenna gains only vary across elevation, thus the path gain is only affected by the elevation AoA and AoD of the MPCs. Note that $\theta$ in Assumption 2 is slightly different from AoD $\theta_{tx,k}$. Both AoA $\theta_{rx,k}$ and AoD $\theta_{tx,k}$ are the angles the MPCs make with the z-axis. The four parameters for each MPC are ultimately affected by objects in the environment and the frequency of operation through the three main mechanisms: (a) reflection from surfaces, (b) diffraction from edges, and (c) transmission through various building materials. Both the building materials and the geometry of the walls can attenuate the MPCs by changing the number of significant reflections, diffractions, and transmissions each path can undergo before it reaches the receiver. We conducted a statistical study on the MPCs generated for an O2I signal propagation scenario relevant to our positioning system. The key results used for further analysis are shown below, with the details of the study presented in Appendix -D.
@@
IV-A ToF based range measurements of the first arriving path
To achieve our ultimate goal of 3D positioning, we are interested in estimating the path delay associated with the diffraction MPCs. Based on the previous section, this is ultimately equivalent to estimating the propagation delay of the first arriving path between the outside anchor and the indoor node. The signal’s propagation delay along the first arriving path will correspond to one of the two diffraction paths and can be estimated using the received signal measurements. Assume that the anchor and node are perfectly time synchronized [^3] This is not a strict requirement. We could also use systems based on round trip time (RTT) as specified in 3GPP and WiFi standards. Now, we transmit a pulse of power $E_{p}$ and pulse shape $p(t)$ of duration $T_{p}$ from the transmitter. Convolving the transmit pulse with the channel defined in eq. (15) we obtain the received signal in the presence of Gaussian noise as